OpenStudy (anonymous):

As you can see, the biker and his friend are having a hard time getting up the hill. They would like to get to the top safely. Two ordered pairs on the hill are given as (6, 3) and (8, 7). Part 1: Find the slope of the hill. Slope (m) = change in y over change in x.

OpenStudy (anonymous):

6 - 8 - 2 1 slope (m) = ------- = ----- = ---- 3 - 7 - 4 2

OpenStudy (anonymous):

Now find the equation using the point slope formula. Point Slope Formula: y – y1 = m (x – x1)